2.1 You are required to evaluate the performance of the selected algorithms from Question 1 by using

the following assumptions in your computation: (15)

ALGORITHM 1

True Positive = 65

True Negative = 70

False Positive = 80

False Negative = 90

ALGORITHM 2

True Positive = 75

True Negative = 80

False Positive = 40

False Negative = 45

Compute the following Machine Learning metrics by showing your working:

Accuracy

Precision

Recall

F1-Measure

2.2 Compare the performance of the selected algorithms from Question 1 based on the Accuracy and

Precision values. (5)

2.3 Contrast the performance of the selected algorithms based on the Recall and F1- Measure. (5)


1
Expert's answer
2021-10-04T23:11:02-0400

ALGORITHM 1

i. Accuracy


Accuracy=correct predictions/total prediction


=(65+70)/(65+80+90+70)


=135/305


=0.442623


ii. Precision


precision=TP/(TP+FP)


=65/(65+80)=65/145


=0.448276

iii. Recall

  recall=TP/(TP+FN)

  =65/(65+90)

  =65/155

  =0.4193548

ALGORITHM 2

i. Accuracy


Accuracy=correct predictions/total prediction


=(75+80)/(75+45+40+80)


=155/240

=0.64583


ii. Precision


precision=TP/(TP+FP)


=75/(75+45)

=75/120

=0.625

iii. Recall

  recall=TP/(TP+FN)

  =75/(75+40)

  =75/115

  =0.652173